### v1

# Electron Magneto-Hydrodynamics in Graphene

Abstract
We consider the hydrodynamic flow of an electron fluid in a channel formed in a two-dimensional electron gas (2DEG) with no-slip boundary conditions. To generate vorticity in the fluid the flow is influenced by an array of micromagnets on the top of 2DEG. We analyse the viscous boundary layer and demonstrate anti-Poiseuille behaviour in this region. Furthermore we predict a longitudinal Hall effect, where a periodic magnetic field generates a voltage modulation in the direction of transport. From the experimental point of view we propose a method for a precise measurements of the properties of different electron fluids. The results are applicable to graphene away from the charge neutrality point and to semiconductors.
I Introduction
It has been over half a century since the theory of viscous electron hydrodynamics was first developed [1] and in recent years viscous effects have been observed in the electron fluid experimentally [2, 3]. Electron transport transitions to the hydrodynamic regime when the momentum-conserving electron-electron scattering length, $l_{ee}$, is the shortest length scale in the system and becomes much shorter than the momentum-nonconserving diffusive scattering length, $l_{\rm mfp}$, and the width of the sample, $W$. In this regime electron-electron interactions dominate and a local thermodynamic equilibrium is established among the electrons, enabling them to flow as a fluid. Since this condition requires samples of very high purity, the theoretical prediction of the hydrodynamic regime in the 1960s [1, 4] could only be realized experimentally after the development of high-mobility semiconductors [5] and graphene [6, 2].
Hydrodynamic electron flow manifests itself in a range of effects that have been observed. While hydrodynamic electron transport was suggested to have been achieved already in 1995 in Ref. [5], this result was immediately challenged by Gurzhi in Ref. [7]: the conditions for hydrodynamics in the bulk material were satisfied, but they were not satisfied for hydrodynamics in a 2DEG. Reliable evidence of the hydrodynamic regime was first presented two decades later in Refs. [8, 9, 10]. Notable consequences of hydrodynamic behavior are the decrease of resistance with increasing temperature known as the Gurzhi effect [5, 11, 12, 13, 14, 15], Poiseuille flow profiles [16, 17, 18], negative non-local resistance and formation of whirlpools [19, 20, 8, 21], the formation of local viscous regions [22], Hall viscosity [14, 23], negative magnetoresistance [24, 25, 13, 26, 10, 27, 28], the violation of the Wiedemann-Franz law [^1] Although recent work in Ref. [29] suggests that the result of a well-known experimental work [30] is not hydrodynamic in nature. [30, 31, 32, 33, 34], anomalous scaling of the resistance with the channel width [31, 9], resonant photoresistance in magnetic fields [35, 36, 37, 38, 27], and a quantum-critical dynamic conductivity [39]. Furthermore, many novel phenomena have been proposed, including the elimination of the Landauer-Sharvin resistance [40], anisotropic fluids [41, 42], dynamo effect in the electron-hole plasma [43] and the prospect of hydrodynamic spin transport [44, 45, 46] inspired by experiments on liquid mercury [47, 48]. In addition to the Poiseuille-like flow expected from electron hydrodynamics, anomalous flow profiles have been predicted in the boundary layer in charge neutral graphene in a uniform perpendicular magnetic field, with this “anti-Poiseuille” flow attributed to recombination effects [49, 50, 51].
Hydrodynamic flow depends qualitatively on boundary conditions, with no-slip boundaries giving rise to the well known Poiseuille flow profile. On the other hand, it is also possible to achieve perfect-slip boundaries by using gates to create a channel in a clean semiconductor. This unique property was used in Ref. [13] for precise measurements of the viscosity in clean GaAs. The measurements indicate deviations from the predictions of the conventional theory of electron-electron scattering and this exciting observation requires more careful studies.
There exists a generic problem in electron hydrodynamics, namely disentangling Ohmic effects, boundary effects, and hydrodynamic effects in experimental data. Viscous dissipation is proportional to the vorticity of the fluid, while the Ohmic dissipation is independent of vorticity. To enhance hydrodynamics one thus needs to enhance the vorticity. This might be achieved by generating turbulent flow of the electron fluid, but the electron fluid is extraordinarily viscous and therefore it is practically impossible to approach turbulence.
In the laminar regime in a uniform magnetic field, vorticity is generated only near the boundaries of the channel in a layer of width $\sim\sqrt{l_{ee}l_{\rm mfp}}$. The flow in this “Knudsen layer” is influenced not only by the viscous fluid effects, but also by the boundary conditions that are device specific and typically not well known. Therefore it is hard to disentangle the unknown boundary effects and unknown viscous effects.
In Ref. [13] the problem was partially resolved by using a gated channel in a GaAs 2DEG that supports perfect-slip boundary conditions. This allowed for relatively accurate measurements of the viscosity of the electron fluid. However, the accuracy of the measurements was limited by Ohmic dissipation which was more than an order of magnitude larger than the viscous one.
In Ref. [52] we suggested a method that involved the use of micromagnets to create wiggling hydrodynamic flow and hence create vorticity throughout the bulk of the sample, not just in the boundary region. This allows for a dramatic enhancement of viscous hydrodynamic dissipation. Hydrodynamics in a finite channel is inherently a boundary problem and in Ref. [52] the problem was solved only for a straight channel in the unique case of perfect-slip boundaries. In the present work we extend the method to a straight channel with arbitrary boundary conditions, either no-slip or finite-slip boundaries. Of course the bulk solution is the same, but the structure of the Knudsen layer depends on the boundary condition and we determine the structure in the present work. Hence, the present work extends the magneto-hydrodynamic method to graphene and also to $\delta$-doped semiconductors.
The structure of this paper is as follows. In Sec. II we briefly describe the solution for the bulk flow through a sinusoidal magnetic field found in Ref. [52]. The difference from Ref. [52] is that we present numerical estimates for graphene instead of GaAs. In Sec. III we study the boundary layer in a straight channel using no-slip boundary condition and characterize the flow near the boundary. We first solve the boundary conditions using perturbation theory and then expand this idea to develop a numerically exact solution. The key results of Sec. III are the width of the boundary layer and the emergence of pseudo anti-Poiseuille flow at high magnetic field strengths. Understanding the width of the boundary layer allows a lower limit to be placed on the sample size in order to use the bulk solution for the “zigzag” flow. In Sec. VI we calculate the magnetic field of periodic ferromagnetic stripes and hence determine parameters of micromagnets necessary for the experimental set up assuming monolayer graphene with experimentally realistic density and mobility. Sec. VII presents our conclusions.
II Bulk Solution for Electron Hydrodynamics in a sinusoidal magnetic field
Consider a two-dimensional channel where $x$ is the longitudinal direction, $y$ the transverse direction and $z$ is perpendicular to the plane. The electron fluid may be controlled through the Lorentz force and thus obeys the modified stationary Navier-Stokes equation at constant density [53],
$\mathbf{v}/\tau-\nu\nabla^{2}\mathbf{v}+\mathbf{v}\cdot{\nabla}\mathbf{v}$ $=-\nabla\Phi/m^{*}+q/m^{*}\mathbf{v}\times\mathbf{B},$
$\nabla\cdot\mathbf{v}$ $=0.$ (1)
Here, $\mathbf{v}$ is the fluid velocity field, $\nu$ is the kinematic viscosity, $\tau$ is the relaxation time related to Ohmic processes (scattering from impurities and phonons), $\Phi$ is the electro-chemical potential, $q$ is the charge of the fluid particle ($q=-e<0$ for electrons), $m^{*}$ is the effective hydrodynamic mass and $\mathbf{B}$ is the magnetic field. For a uniform field, it is known that this equation predicts the usual Hall effect, whereas recently it has been realized that non-uniform fields have remarkable consequences such as a local suppression of flow [54] and vortex layers in the bulk [52]. The electron-electron scattering length, $l_{ee}$, and the momentum diffusive scattering length, $l_{\rm mfp}$, are defined by the following relations
$\nu=\frac{l_{ee}v_{F}}{4}$
$\tau=\frac{l_{\rm mfp}}{v_{F}},$ (2)
where $v_{F}$ is the Fermi velocity at zero temperature.
A superlattice of micromagnets modulated in the $x$ direction generates a sinusoidal magnetic field along the $z$ direction in the plane of the 2DEG,
$B_{z}=B_{0}\sin(gx)\ ,$ (3)
where $g=2\pi/a$ and $a$ is the separation between adjacent micromagnets. As previously mentioned, the electron fluid is very viscous, so the “Bernoulli term”, $\mathbf{v}\cdot{\nabla}\mathbf{v}$, in the left hand side of Eq. (II) can be neglected. Hence, we obtain a linear Stokes equation that can be solved using the stream function $\psi$ defined by $v_{x}=\partial_{y}\psi$ and $v_{y}=-\partial_{x}\psi$. For an infinitely long and infinitely wide channel with fixed averaged current density $\mathbf{J}=qnv_{0}\mathbf{\hat{x}}$ the solution is
$\psi_{\infty}=v_{0}\left(y+\frac{\omega\tau}{g(1+D^{2}g^{2})}\cos(gx)\right),$ (4)
where $D=\sqrt{\nu\tau}=\frac{1}{2}\sqrt{l_{ee}l_{\rm mfp}}$ is the characteristic viscous length and $\omega=|e|B_{0}/m^{*}$ is the cyclotron frequency. The viscous and Ohmic dissipation rates follow directly from Eq. (4), were $L$ and $W$ are the length and the width of the channel, respectively,
$\dot{E}_{\Omega}$ $=$ $\frac{nm^{*}}{\tau}\int d^{2}r\>v^{2},$
$\dot{E}_{\nu}$ $=$ $nm^{*}\nu\int d^{2}r\>\mathbf{v}\cdot\nabla^{2}\mathbf{v}$ (5)
For the bulk solution (4) this gives
$\dot{E}_{\Omega}$ $=$ $\frac{nm^{*}v_{0}^{2}}{\tau}\ \left(1+\frac{\omega^{2}\tau^{2}}{2(1+D^{2}g^{2})^{2}}\right)LW,$
$\dot{E}_{\nu}$ $=$ $\frac{nm^{*}v_{0}^{2}}{\tau}\ \frac{g^{2}D^{2}\omega^{2}\tau^{2}}{2(1+D^{2}g^{2})^{2}}LW.$ (6)
The coefficient $nm^{*}v_{0}^{2}LW/\tau$ can be identified as the dissipation due to the Drude resistance in a conventional conductor with diffusive transport. The ratio of viscous to Ohmic dissipation is dependent on the dimensionless parameters $\omega\tau$ and $gD$ which, for fixed $\omega\tau$, may be shown to be maximized at $gD=[1+(\omega\tau)^{2}/2]^{1/4}$.
So far we have not explicitly solved for the electro-chemical potential in the Stokes equation (II). Substituting $\mathbf{v}$ calculated from the stream function (4), it is easy to derive an expression for the electro-chemical potential, representing the external force on each fluid particle,
$\partial_{x}\Phi=-\frac{m^{*}v_{0}}{\tau}\left(1+\frac{\omega^{2}\tau^{2}\sin^{2}(gx)}{1+D^{2}g^{2}}\right),$
$\partial_{y}\Phi=0.$ (7)
The electro-chemical potential does not vary in $y$ in an infinite sample as there is translational symmetry in the $y$ direction, whereas the periodic structure of the magnetic field in the $x$ direction gives rise to a modulated electro-chemical potential as a function of $x$. Furthermore, it is clear that the rate of work done by the external force $\dot{E}=-\int d^{2}r\>\mathbf{v}\cdot\nabla\Phi$ matches the total dissipation $\dot{E}_{\Omega}+\dot{E}_{\nu}$ obtained in Eq. (II).
As the electron number density is approximately constant throughout the sample, the thermodynamic pressure gradients vanish. Therefore, the electrochemical potential gradients are predominantly electric in nature, and we can write the voltage as $\varphi=-\Phi/|e|$,
$\varphi=\frac{m^{*}v_{0}}{|e|\tau}\left(x+\frac{\omega^{2}\tau^{2}}{2(1+D^{2}g^{2})}\left(x-\frac{\sin(2gx)}{2g}\right)\right).$ (8)
From Eq. (8) it is clear that the electric field varies with $x$. As this effect arises from the presence of a periodic magnetic field and the modulation in voltage is in the same direction as transport it is appropriate to label this a “longitudinal Hall effect”.
[…]
$n=10^{12}\>\text{cm}^{-2}$
$T=220\>\text{K}$
$l_{\rm mfp}=1.9\>\mu\text{m}$
$l_{ee}=0.24\>\mu\text{m}$
$\tau=1.9\>\text{ps}$ (9)
[…]
With these parameters and assuming $v_{0}=100$ m/s, corresponding to $I=0.3$ $\mu$A in a narrow $W=2$ $\mu$m channel from Ref. [8], the oscillation amplitude of the longitudinal Hall effect is approximately $0.3$ $\mu$V.
III Boundary layer for a straight semi-infinite channel with no-slip boundary condition
III.1 Boundary layer flow in the perturbative limit $\omega\tau\ll 1$
Consider a semi-infinite straight channel with a no-slip boundary at $y=0$ and fluid flow in the region $y>0$. In conjunction with the Stokes equation (II), the stream function must satisfy the no-penetration and no-slip boundary conditions,
$v_{y}|_{y=0}=-\partial_{x}\psi|_{y=0}=0,\;v_{x}|_{y=0}=\partial_{y}\psi|_{y=0}=0.$ (10)
The no-penetration condition ensures that no fluid can cross the boundary whereas the no-slip condition ensures that the fluid has zero tangential velocity at the boundary surface. Furthermore, the stream function must decay to the bulk solution (4) far from the boundary. We look for a solution as an expansion in powers of $\omega\tau\ll 1$, i.e., $\psi=\sum_{n=0}^{\infty}(\omega\tau)^{n}\psi^{(n)}$. First we solve for the zero-order solution, corresponding to flow in the absence of a magnetic field. Solving the Stokes equation for $\omega\tau=0$ using the boundary conditions leads to the Poiseuille solution
$\psi^{(0)}=v_{0}\left(y+De^{-y/D}\right).$ (11)
Next, introducing the magnetic field as a perturbation we derive the first-order correction in the solution, $\psi=\psi^{(0)}+\omega\tau\psi^{(1)}$,
$\psi^{(1)}=v_{0}\cos(gx)\left(\frac{1}{g(1+g^{2}D^{2})}+Ae^{-y/D}\right.$
$+\left.C_{1}e^{-gy}+C_{2}e^{-\sqrt{g^{2}+\frac{1}{D^{2}}}y}\right),$ (12)
where the coefficients $A$, $C_{1}$ and $C_{2}$ are defined as
$A$ $=\frac{1}{g(1-g^{2}D^{2})},$
$C_{1}$ $=\frac{2(g^{2}D^{2}+1)^{-\frac{1}{2}}-1}{g(1-g^{2}D^{2})(gD-\sqrt{g^{2}D^{2}+1})},$
$C_{2}$ $=\frac{1-gD}{g(1+gD)(1+g^{2}D^{2})(gD-\sqrt{g^{2}D^{2}+1})}.$
Note that in the case of $gD\to 1$ the coefficients $A,C_{1}\to\infty$. However, it is easy to check that the divergences compensate each other such that the solution remains finite.
To maximize the viscous dissipation we are interested in the case $gD>1$. Since the slowest decaying boundary term in Eq. (III.1) decays as $e^{-y/D}$, the characteristic viscous length $D$ sets the width of the boundary layer.
III.2 Exact numerical solution for the boundary layer flow
While the perturbative solution for the boundary layer presented in the previous subsection is illuminating, for practical purposes we must find the solution at $\omega\tau>1$. This can only be done numerically. To do so we reduce the Stokes equation (II) to an infinite series of linear algebraic equations. We then numerically solve these equations by truncating them at a suitable order.
Figure 1: Streamlines in a semi-infinite channel with a no-slip boundary at $y=0$. The parameters are chosen as $gD=1.16$, $\omega\tau=0.3,1.2,4.5$. Solid lines correspond to the numerically exact solution described in Sec. III.2 and dashed lines correspond to the solution given by Eq. (4) in the infinite device.
We consider a semi-infinite channel $y\geq 0$ and use length units where $D=\sqrt{\nu\tau}=1$ and velocity units where $v_{0}=1$. As the magnetic field is periodic in the $x$ direction, so is the stream function $\psi$, which can thus be presented as the following Fourier series,
$\psi=y+e^{-y}+\frac{\omega\tau}{g(g^{2}+1)}\cos(gx)$
$+\sum_{m=-\infty}^{\infty}e^{igmx}G_{m}(y)\ .$ (13)
Here $m$ is an integer and $G_{m}(y)$ are unknown functions that must decay with increasing $y$ as the solution tends to the bulk solution far from the boundary. Thus, using the perturbative solution as a starting point we consider solutions of the form
$G_{m}(y)=f_{m}e^{-y}+\sum_{a}A_{m,a}e^{-\lambda_{a}y},$ (14)
where $f_{m}$ and $A_{m,a}$ are complex coefficients. The index $a=1,2,3...$ enumerates “eigenvalues” $\lambda_{a}$, which can be real or complex. The real parts of $\lambda_{a}$ are positive by design of Eq. (14). In principle the index $a$ can be arbitrarily large, but due to the truncation it is never larger than $\sim 10$. Since the “Bernoulli term” is neglected, the Stokes equation (II) is a linear equation. Let us consider the Poiseuille solution without magnetic field,
$\psi_{0}=y+e^{-y}\ ,$ (15)
as a base solution. Now, in a magnetic field the stream function has a driven part correction, $\delta\psi_{d}$, driven by $\psi_{0}$, and a free part correction, $\delta\psi_{f}$,
$\psi$ $=\psi_{0}+\delta\psi_{d}+\delta\psi_{f},$
$\delta\psi_{d}$ $=\frac{\omega\tau}{g(g^{2}+1)}\cos(gx)+\sum_{m=-\infty}^{\infty}e^{igmx}f_{m}e^{-y},$
$\delta\psi_{f}$ $=\sum_{m=-\infty}^{\infty}e^{igmx}\sum_{a}A_{m,a}e^{-\lambda_{a}y}$ (16)
The first term in $\delta\psi_{d}$ is driven by the first term in Eq. (15) and the second term in $\delta\psi_{d}$ is driven by the second term in Eq. (15). Hence, the function $G_{m}(y)$ in Eq. (14) is a combination of a driven part and a free part. For a linear differential equation the driven part is fully determined by the equation, but the coefficients in the free part are determined by the boundary conditions.
It is convenient to define the following differential operators
${\hat{D}}$ $\equiv-\nabla^{2}+\nabla^{4},$
${\hat{d}}$ $\equiv\omega\tau g\cos(gx)\partial_{y}.$ (17)
Using these notations the Stokes equation can be rewritten as
$(\hat{D}-{\hat{d}})\delta\psi_{d}$ $=\hat{d}\psi_{0},$
$(\hat{D}-\hat{d})\delta\psi_{f}$ $=0$ (18)
Using the representation (III.2) the driven equation is transformed to an infinite set of linear algebraic equations for the coefficients $f_{m}$
$(g^{2}m^{2})(g^{2}m^{2}-1)f_{m}+\frac{\omega\tau g}{2}(f_{m+1}+f_{m-1})\\ =-\frac{\omega\tau g}{2}(\delta_{m,1}+\delta_{m,-1}).$ (19)
The coefficients in this equation are real and even in $m$. Therefore, the coefficients $f_{m}$ are also real and even in $m$. For this reason only equations for non-negative $m$ are independent. A numerical solution of Eq. (19) is straightforward, with truncation beyond $m>N_{m}$, resulting in $N_{m}+1$ independent equations. The second equation in Eq.(III.2), i.e., the free equation, is transformed to an infinite set of linear algebraic equations for the coefficients $A_{m}$,
$(g^{2}m^{2}-\lambda_{a}^{2})(1+g^{2}m^{2}-\lambda_{a}^{2})A_{m,a}$
$+\frac{\omega\tau g}{2}\lambda_{a}(A_{m+1,a}+A_{m-1,a})=0,$ (20)
noting that $\lambda_{a}$ is independent of $m$. Again, only equations with non-negative m are independent and we truncate $m$ such that $m\leq N_{m}$. Hence the number of independent equations is again $N_{m}+1$. The right hand side of Eq. (III.2) is zero, so for a nontrivial solution the determinant of the matrix must be equal to zero. Hence, this is an eigenvalue problem but it is not a standard linear eigenvalue problem. To solve the eigenvalue problem we calculate the determinant as a polynomial in $\lambda$ and numerically determine the roots of the polynomial. The polynomial order is $4(N_{m}+1)$. After we discard roots with non-positive real part we are left with $12N_{m}+$ 1 roots. Interestingly, at a sufficiently large $\omega\tau$ there are complex roots. The coefficients in the polynomial are real, so complex roots always come in complex conjugate pairs. We enumerate $12N_{m}+$ 1 roots by the index $a$. For any found $\lambda_{a}$ we solve Eqs. (III.2) and hence find the coefficients $A_{m,a}$, numerically determined to arbitrary normalization. We now write $A_{m,a}=\alpha_{a}A_{m,a}^{(n)}$, where $A_{m,a}^{(n)}$ is normalized such that $\sum_{m}|A_{m,a}^{(n)}|^{2}=1$, and $\alpha_{a}$ is an arbitrary coefficient.
The boundary conditions (10) provide a set of equations to determine the $12N_{m}+$ 1 coefficients $\alpha_{a}$
$0=v_{y}|_{y=0}\Longrightarrow$
$m(f_{m}+\sum_{a}\alpha_{a}A_{m,a}^{(n)})+\frac{\omega\tau}{2g(1+g^{2})}(\delta_{m,1}-\delta_{m,-1})=0,$
$0=v_{x}|_{y=0}\Longrightarrow$
$f_{m}+\sum_{a}\lambda_{a}\alpha_{a}A_{m,a}^{(n)}=0.$ (21)
The first of these equations is nontrivial at $m=1,2...N_{m}$ and the second equation is nontrivial at $m=0,1...N_{m}$, so altogether we have $12N_{m}+$ 1 equations to determine $12N_{m}+$ 1 unknown coefficients $\alpha_{a}$. Thus the boundary layer problem is solved.
The boundary layer solution considered in this Section is valid for the no-slip boundary condition. However, the method can be easily extended to the case of an arbitrary slip length, considering that only Eqs. (III.2) depend on the slip length.
The boundary layer solution is extended to finite width channels in App. A. After the method is developed it is instructive to present plots for specific cases. For the numerical solution we set $gD=1.16$, which corresponds to the device described at the end of Sec. II. We truncate beyond $N_{m}=5$, which provides more than sufficient accuracy for the values of $\omega\tau$ we present. Plots of the streamlines obtained for $\omega\tau=0.3$, $1.2$ and $4.5$ are presented in Fig. 1. These values of $\omega\tau$ were selected as $\omega\tau=0.3$ has only real eigenvalues $\lambda_{a}$ and corresponds to the perturbative regime, $\omega\tau=1.2$ has a complex conjugate pair of eigenvalues $\lambda_{a}$ and corresponds to the device described in Eq. (II) and $\omega\tau=4.5$ has two complex conjugate pairs of eigenvalues $\lambda_{a}$. As the magnetic field increases ($\omega\tau$ increases) both the magnitude of the boundary effects and the width of the boundary layer increase, but in the cases considered the effective width of the boundary layer does not extend beyond $y/D\approx 4$. Thus, considering symmetry, it is appropriate to use the bulk solution in devices with width $W\gg 8D$.
IV Numerical estimates, anti-Poiseuille flow
Following the numerical calculation of the stream function, we may calculate the the viscous dissipation rate and the Ohmic dissipation rate given by Eqs. (II). The most important parameter is the ratio $\dot{E}_{\nu}/\dot{E}_{\Omega}$, which is dependent on the device considered, so to provide calculated values we take the device with parameters presented in Eq. (II). This corresponds to $D=0.34\>\mu\text{m}$. We also take the lattice spacing $a=1.9\>\mu\text{m}$ as discussed in the end of Section II, resulting in $gD=1.16$. The ratio $\dot{E}_{\nu}/\dot{E}_{\Omega}$ is plotted in Fig. 2 versus the channel width $W$, for two values of the magnetic field amplitude, $B_{0}=75$ mT,$150$ mT, corresponding to $\omega\tau=1.2$, $2.4$ respectively. In these plots we assume that the viscosity is independent of the magnetic field, $\nu(B)=\nu(0)$. Asymptotic values of the ratio $\dot{E}_{\nu}/\dot{E}_{\Omega}$ at $W\to\infty$ are $0.16$ ($B_{0}=75$ mT) and $0.48$ ($B_{0}=150$ mT).
Figure 2: Ohmic (blue line) and viscous (red line) dissipation as a function of the width of the channel normalized to the $\dot{E}_{0}$ which is due to Drude resistance only, i.e., what is expected in the absence of B-field and hydrodynamics ($\dot{E}_{0}/LW=nm^{*}v_{0}^{2}/\tau$). The value $\omega\tau=1.2$ corresponds to $B_{0}=75$ mT and $\omega\tau=2.4$ corresponds to $B_{0}=150$ mT. In these plots we assume that the viscosity is independent of the magnetic field, $\nu(B)=\nu(0)$. The plots correspond to the parameter set (II), where $gD=1.16$.
However, the viscosity depends on magnetic field, with possible dependence [10]
$\nu\propto\frac{1}{1+(B/B^{*})^{2}}\;,\;B^{*}=\frac{p_{F}}{2|e|l_{ee}}$ (22)
For the device parameters (II) the characteristic magnetic field is $B^{*}\approx 240\>\text{mT}$. Therefore, for $B_{0}=150$ mT the approximation $\nu(B)=\nu(0)$ is questionable, but the possible correction does not exceed 10-20%. The viscous dissipation comes from the regions of maximum vorticity and in this setup the vorticity is maximum at the regions where $B\to 0$, so the true correction for $B_{0}=150$ mT is lower than that given simply by Eq. (22). It is not a problem to solve the Navier-Stokes equation with position dependent viscosity, but it is not necessary at this stage.
Interestingly, the fluid flow in the Knudsen layer shows a sort of anti-Poiseuille behaviour. To illustrate this, we plot in Fig. 3 the longitudinal velocity $v_{x}$ versus $y$ (i.e., across the channel). The velocity $v_{x}$ is a function of both $x$ and $y$ and here we present plots at $x=0$. At a sufficiently large magnetic field, $\omega\tau\geq 2.4$, there is a clear anti-Poiseuille profile near the left boundary of the channel. In this regime the peak velocity is not $v_{0}$ in the bulk, but instead the longitudinal velocity has a maximum in the boundary layer before decaying towards $v_{0}$ in the bulk. There is a conventional Poiseuille behaviour at the right boundary of the channel, although at high magnetic field a small amount of counterflow develops due to the development of vortices in the boundary layer. If we shift $x=0\to x=a/2$ the plots are effectively reflected about the centre of the channel: the anti-Poiseuille profile is located at the right boundary and the Poiseuille profile is at the left boundary. Mathematically, the anti-Poiseuille profile is related to the complex eigenvalues $\lambda_{a}$. This behaviour is qualitatively similar to the anti-Poiseuille flow predicted for narrow graphene channels near charge neutrality in the presence of a uniform magnetic field perpendicular to the plane [49, 50, 51]. Near charge neutrality this effect is largely attributed to electron-hole recombination. In our case the physical origin of the anti-Poiseuille profile is different as we do not have any holes. While we consider no-slip boundaries, a similar anti-Poiseuille profile exists also for boundary conditions with a finite slip length. The location of the anti-Poiseuille maximum may be used to extract the slip length in a sample with finite-slip boundaries.
Figure 3: Longitudinal velocity profile at $x/D=0$ with $gD=1.16$ for various values of $\omega\tau$. At large magnetic fields anti-Poiseuille flow is observed and a small counterflow is observed at the opposite boundary due to the formation of a viscous vortex.
V Finite-elements simulations
To complement our numerical results and to extend them towards finite systems, we also performed numerical simulations using the finite-element method (FEM). The first step is to bring Eq. (II) to the form of a weak differential equation by multiplying each of the equation with a test function and integrating over all space. We again neglect the Bernoulli term because we assume small flow velocity and assume the fluid to be incompressible. For the finite-element analysis, we use the space of Lagrange polynomials of second order as the function space for the velocity field and the space of Lagrange polynomials of first order as the function space for the potential. Together these form a Taylor-Hood element appropriate for the numerical solution of Navier-Stokes equations. We have numerically implemented the problem using the FEniCS package [55].
Using FEM, we consider a finite system with a rectangular geometry with width $W$ in $y$ direction and length $L$ in $x$ direction. To simulate a finite array of micromagnets, we consider a flow driven by an applied electric potential difference, so we assume Dirichlet boundary conditions for the voltage $\varphi(x=-L/2,y)=\varphi_{L}$ and $\varphi(x=L/2,y)=\varphi_{R}$ at inlet and outlet. In addition to this, we assume that $v_{y}(x=\pm L/2,y)=0$ at both inlet and outlet. Finally, we impose no-slip boundary conditions at the walls, i.e., $\mathbf{v}(x,y=\pm W/2)=0$.
Figure 4: FEM results for the absolute value $|\mathbf{v}|$ of the velocity in the channel for $gD=0.5$. Note that this color plot should not be confused with lines of flow, which are shown in Fig. 1 The velocity is measured in units convenient for numerical solution and the system contains six magnetic strips. The region of high velocity, anti-Poiseuille flow, is shifted by half a period at opposite boundaries.
Figure 4 shows that the results of these simulations qualitatively reproduce all features found for the infinite system. At strong enough magnetic fields, an anti-Poiseuille flow emerges near the boundary. Moreover, at strong fields, a small counterflow becomes visible at the opposite boundary.
VI Magnetic field of periodic ferromagnetic stripes
The setup considered in this paper is similar to the setup previously proposed for GaAs heterostructure in Ref. [52], but adapted to suit monolayer graphene with parameters described in Eq.(II). A superlattice of rectangular magnetic bars oriented perpendicular to the longitudinal axis of the device is placed above the 2DEG as seen in the left panel of Fig. 5.
Figure 5: Left: top view of a periodic array of magnetic bars perpendicular to the flow direction shown by the red arrow. The period of the array is $a$ and the width of the bar is $d$. The bars are magnetized in parallel to the longitudinal axis of the channel. The direction of the magnetization is shown by the blue arrows. Right: Side view of the device. The vertical distance from the bar magnets to the 2DEG is $z$ and the thickness of the rectangular bar is $\delta z$.
The bars are magnetized along the longitudinal axis of the channel. The bar thickness is $\delta z$ and the distance from the base of the bar to the 2DEG is $z$, shown in the right panel of Fig. 5. The period of modulation is $a=1.9$ $\mu$m as found in Sec. II. We tune the remaining free parameters $d,z,\delta z$ such that the $z$ component of the magnetic field in the 2DEG is approximately $B_{0}\sin(2\pi x/a)$ with $B_{0}\approx 75$ mT. A stronger magnetic field may be also achieved.
Assuming the magnetization is uniform throughout the volume of each bar, and the bars’ extent in the transverse ($y$) direction is wider than the width $W$ of the sample, we calculate the magnetic field in the plane of the 2DEG as a function of longitudinal coordinate $x$ from the Biot-Savart law as
$\frac{B_{z}(x)}{B_{m}}\bigg{|}_{\rm 2DEG}$ $=-\sum_{n>1}\frac{\sin\left(gnd/2\right)}{\pi n}\sin(gnx)Z_{n},$
$Z_{n}(z,\delta z)$ $=e^{-gnz}(1-e^{-gn\delta z}),$ (23)
where $B_{m}$ is the saturation field, and $x$ is measured from the center of one of the magnets. For example, at the edge of the magnet $x=d/2$, the harmonics of the field in the 2DEG plane is $\propto\sin^{2}(gnd/2)$. Therefore, $d=a/2=\pi/g$ maximizes the first harmonic and suppresses the second harmonic. We numerically calculate $B$ from Eq. (VI) and further confirm that $d=a/2=0.93\>\mu$m is optimal.
In order to achieve a field with magnitude $B_{0}=75$ mT we consider a NiFe alloy that has saturation magnetic field $1500$ mT. [56] Through concurrently tuning $\delta z$ and $z$ we find that we may generate the desired magnetic field profile with $\delta z=0.15$ $\mu$m and $z=0.30$ $\mu$m. Plots of $B_{z}$ versus $x$ and $B_{x}$ versus $x$ are presented in Fig. 6.
Figure 6: Magnetic field in the 2DEG plane for NiFe bars. Left panel: $B_{z}$ versus $x$. Right panel: $B_{x}$ versus $x$. The parameters are $\delta z=0.15\>\mu\mathrm{m},\ z=0.30\>\mu\mathrm{m}$.
The $x$ component of the magnetic field is not relevant, but we present it for completeness. The $x$ dependence of $B_{z}$ is not quite sinusoidal, however this is not a big issue as using the finite element method one can also solve the Navier-Stokes equation for this magnetic field profile.
VII Conclusions
We develop a theory for the no-slip boundary layer in the magnetohydrodynamic flow of an electron fluid in a straight channel. The fluid flow is modified by an array of micromagnets on the top of two-dimensional electron gas. The micromagnets create a wiggling flow of the fluid, which dramatically enhances hydrodynamic effects in the fluid. To be specific, the analysis was performed for no-slip boundary conditions, but the method can be extended to boundary conditions with arbitrary slip length. We demonstrate that the fluid velocity profile in the boundary layer manifests an anti-Poiseuille behaviour. In addition, we show that the micromagnet array creates a longitudinal Hall effect in the sample. From the experimental point of view we propose a method for a precise measurements of the properties of different electron fluids. The results are applicable to graphene away from the charge neutrality point and also to semiconductors.
Acknowledgements
We acknowledge important discussions with Alexander Hamilton, Yonatan Ashlea Alava, Oleh Klochan, Daisy Wang and Zeb Krix. This work was supported by the Australian Research Council Centre of Excellence in Future Low- Energy Electronics Technologies (CE170100039).
Appendix A The method for exact solution for a finite width channel
Here we expand the method of Sec.III to a finite channel with no slip boundaries at $y=\pm W/2$. Here we use units $D=v_{0}=1$. As before we split the stream function into the zero field term which corresponds to Poiseuille flow, $\psi_{0}$, the driven magnetic field correction term, $\delta\psi_{f}$ and the free magnetic field correction term $\delta\psi_{d}$.
$\psi$ $=\psi_{0}+\delta\psi_{d}+\delta\psi_{f},$
$\psi_{0}$ $=y-\frac{\sinh{y}}{\cosh{W/2}}$
$\delta\psi_{d}$ $=\frac{\omega\tau}{g(g^{2}+1)}\cos(gx)+\sum_{m=-\infty}^{\infty}e^{igmx}\left(f_{m}e^{-y}+h_{m}e^{y}\right),$
$\delta\psi_{f}$ $=\sum_{m=-\infty}^{\infty}e^{igmx}\sum_{a}\left(A_{m,a}e^{-\lambda_{a}y}+C_{m,a}e^{\lambda_{a}y}\right)$ (24)
Note that we absorb the constant exponential terms into the unknown coefficients. The stream function solves the Stokes’ equation as in Eq.(III.2). From the driven equation we obtain a set of equations for coefficients $f_{m}$ and $h_{m}$.
$(g^{2}m^{2})(g^{2}m^{2}-1)f_{m}+\frac{\omega\tau g}{2}(f_{m+1}+f_{m-1})$ (25)
$=-\frac{\omega\tau g}{4\cosh{W/2}}(\delta_{m,1}+\delta_{m,-1})$
$(g^{2}m^{2})(g^{2}m^{2}-1)h_{m}-\frac{\omega\tau g}{2}(h_{m+1}+h_{m-1})$
$=-\frac{\omega\tau g}{4\cosh{W/2}}(\delta_{m,1}+\delta_{m,-1})$
As the coefficients are real and even in $m$, $f_{m}$ and $h_{m}$ are also real and even in $m$. The free equation gives eigenequations for eigenvalues $\lambda_{a}$ and eigenvectors $A_{m,a}$ and $C_{m,a}$.
$(g^{2}m^{2}-\lambda_{a}^{2})(1+g^{2}m^{2}-\lambda_{a}^{2}))A_{m,a}$
$+\frac{\omega\tau g}{2}\lambda_{a}(A_{m+1}+A_{m-1})=0$
$(g^{2}m^{2}-\lambda_{a}^{2})(1+g^{2}m^{2}-\lambda_{a}^{2}))C_{m,a}$
$-\frac{\omega\tau g}{2}\lambda_{a}(C_{m+1}+C_{m-1})=0$ (26)
Note that the determinant of the $A$-equation is identical to that of the $C$-equation. Hence, there is only one set of eigenvalues $\lambda_{a}$ which we find following the method described in the main text. The eigenvectors $A_{m,a}$ and $C_{m,a}$ are different and we find them solving Eqs. (A). The eigenvectors are determined up to some arbitrary normalization. Similarly to the approach in Sec.III, we define coefficients $\alpha_{a}$ and $\gamma_{a}$ and write $A_{m,a}=\alpha_{a}A_{m,a}^{(n)}$ and $C_{m,a}=\gamma_{a}C_{m,a}^{(n)}$, where $\sum_{m}\left|A_{m,a}^{(n)}\right|^{2}=1$ and $\sum_{m}\left|C_{m,a}^{(n)}\right|^{2}=1$, and $\alpha_{a}$ and $\gamma_{a}$ are arbitrary coefficients. These coefficients are determined by considering the boundary conditions at $y=\pm\frac{W}{2}$.
$0=v_{y}|_{y=\pm W/2}\to$
$m(f_{m}e^{\mp W/2}+h_{m}e^{\pm W/2}+\sum_{a}(\alpha_{a}A_{m,a}^{(n)})e^{\mp\lambda_{a}W/2}$
$+\gamma_{a}C_{m,a}^{(n)}e^{\pm\lambda_{a}W/2}))+\frac{\omega\tau}{2g(1+g^{2})}(\delta_{m,1}-\delta_{m,-1})=0$
$0=v_{x}|_{y=\pm W/2}\to$
$\sum_{a}\lambda_{a}(\alpha_{a}A_{m,a}^{(n)}e^{\mp\lambda_{a}W/2}-\gamma_{a}C_{m,a}^{(n)})$
$+f_{m}e^{\mp W/2}-h_{m}e^{\pm W/2}+e^{\pm\lambda_{a}W/2}=0$ (27)
Upon solving Eq. (A) and obtaining $\alpha_{a}$ and $\gamma_{a}$ the stream function is entirely characterized and velocity and dissipation profiles may be calculated.